OpenStudy (anonymous):

There is a unique real root to the following. Solve with pen and paper.\[x^7+7x^3-7x^2+7x+1=0\]

OpenStudy (anonymous):

Newton-Ralphson method..? Or trial evaluation, eventually zooming in on a fair interval?

OpenStudy (anonymous):

Either way, \(x\approx -0.125.\)

OpenStudy (anonymous):

clearly no real solution for \(x\ge0\) and \(x\le −\frac{1}{7}\) so we must work on \[−\frac{1}{7}<x<0\] for this interval \(x^7≈0\) and also \(x^3≈0\) solving \(−7x^2+7x+1=0\) gives \(x≈−0.126\)

OpenStudy (anonymous):

actually from quadratic \[x=\frac{7-\sqrt{77}}{14}≈-0.126\]